Session: May 1998 (session 2)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (T. Venturi at Thu Jun 4 19:41:46 METDST 1998) Success
No apparent problems were spotted

Westerbork (Tony Foley at Fri Jun 5 09:09:15 METDST 1998) Minor failures
Some problems with the decoder, but otherwise looks OK.

Onsala (A. Polatidis at Mon Jun 8 17:29:58 METDST 1998) Success
No known problems. Tape will be sent on June 9

Noto (C.Trigilio at Mon Jun 8 17:36:59 METDST 1998) Success

Jodrell (Alastair Gunn at Tue Jun 9 03:00:19 METDST 1998) Success
No known problems.

Calibration (Jean\-Francois Desmurs at Wed Jun 10 15:54:06 METDST 1998) Minor failures
The ANTAB formatted file for this experiment can be found on VLBEER in the ? directory under the name fr002.antab. Rem: Jb VC07 strange values. Read the comments in the file to get further information.

Effelsberg (dgraham at Fri Jun 12 14:51:01 METDST 1998) Minor failures
radar interference at 1403. Some problems with overflow in tsys values